OpenStudy (anonymous):

The points (-10,-4) and (5,5) are end points of the diameter of a circle. Find the length of the radius of the circle. Help please?

jigglypuff314 (jigglypuff314):

\[distance formula: \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]when given points (x1, y1) and (x2, y2)

jigglypuff314 (jigglypuff314):

so do\[\sqrt{(5+ 10)^{2} + (5+4)^{2}} =\]then divide what you get by 2

OpenStudy (anonymous):

48?

jigglypuff314 (jigglypuff314):

mmm I got √ (15^2 + 9^2) = 17.493 divided by 2 = 8.746

OpenStudy (anonymous):

what did you do with the (15^2 + 9^2)?

jigglypuff314 (jigglypuff314):

square rooted it

OpenStudy (anonymous):

8.75 .

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

how about this one? (4, -4) and 7,5)

jigglypuff314 (jigglypuff314):

to find the distance between two points \[\sqrt{(7-4)^{2} + (5+4)^{2}}\]

OpenStudy (anonymous):

6.4807?

jigglypuff314 (jigglypuff314):

mmm not quite :) √( 3^2 + 9^2) = ?

OpenStudy (anonymous):

9.4868?

jigglypuff314 (jigglypuff314):

that's what I got :)